The Treasury bill rate is 4%, and the expected return on the market portfolio is 11%. According to the capital asset pricing model:


a. What is the risk premium on the market?
b. What is the required return on an investment with a beta of 1.6? (Do not round intermediate calculations. Enter your answer as a percent rounded to 1 decimal place.)
c. If an investment with a beta of 0.8 offers an expected return of 8.6%, does it have a positive or negative NPV?
d. If the market expects a return of 11.0% from stock X, what is its beta? (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

a.

Risk Premium = Expected Rate on Market - Risk Free Rate

Risk Premium = 0.11 - 0.04

Risk Premium = 7.00%

b.

Using CAPM Model,

Required Rate = Rf + Beta(Risk Premium)

Required Rate = 0.04 + 1.6(0.07)

Required Rate = 15.20%

c.

Required Rate = 0.04 + 0.80(0.07)

Required Rate = 9.6%

But Expected Rate = 8.6%

So, Investment has negative NPV.

d.

0.11 = 0.04 + Beta(0.07)

Beta = 1